OpenStudy (anonymous):

If X is a random variable with moment generating function phi(t) = exp{ λ(e^t − 1)}, find the mean and variance of X.

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

Expected value is just the \(1\)st moment.

OpenStudy (anonymous):

what will the mean be? @wio :(

OpenStudy (anonymous):

The \(n\)th moment is given by the \(n\)th derivative of the moment generating function.

OpenStudy (anonymous):

\[ \text{E}[X] = \phi'(0) \]\[ \text{Var}[X]=\text{E}[(X-\mu)^2] = \text{E}[X^2]-(\text{E}[X])^2 = \phi''(0)-(\phi'(0))^2 \]

OpenStudy (anonymous):

@wio it seems the answer is 0!

OpenStudy (anonymous):

Does that make sense in this case?

OpenStudy (anonymous):

wait, after differentiating, we have put in the value 0 in t.

OpenStudy (anonymous):

I'm getting \(\phi'(0) = \lambda \)

OpenStudy (anonymous):

isn't differentiating that hard? if you do not mind, could you show me in details? Sorry! It would be of great help @wio

OpenStudy (anonymous):

\[ \frac d{dt}\color{blue} \exp[ \color{red}{\lambda(e^t − 1)}] = \color{blue} \exp[\lambda(e^t − 1)]\cdot \color{red}{\lambda e^t} \]

OpenStudy (anonymous):

\[ \frac d{dt} \color{blue} {\exp[\lambda(e^t − 1)]}\cdot \color{red}{\lambda e^t} = \color{blue} {\exp[\lambda(e^t − 1)] \cdot \lambda e^t} \cdot \color{red}{\lambda e^t} + \color{blue} {\exp[\lambda(e^t − 1)]}\cdot\color{red}{\lambda e^t} \]

OpenStudy (anonymous):

For the first one, I use chain rule. Red for inner function blue for outer. For the second one, I use product rule.

OpenStudy (anonymous):

So I'm getting \[ E[X] = \lambda\quad E[X^2]=\lambda^2+\lambda \]

OpenStudy (anonymous):

Heh, I knew something was up when the variance and mean were equal... This is the MGF of the Poisson distribution. http://en.wikipedia.org/wiki/Poisson_distribution

OpenStudy (anonymous):

oh my. this is huge! Thanks a lot :) Var could be easily found by the formular. if we get E[X], den finding VAR wouldn't be tough. Hope I am right?

OpenStudy (anonymous):

I already told you how to get variance...

OpenStudy (anonymous):

right!